If a card is drawn from a deck of playing cards, then what is the probability of getting

a) a red card or a face card?

b) A diamond or a number below 6?

c) a black card or a spade?


1
Expert's answer
2021-10-06T12:15:13-0400

a)

"P=\frac{1}{2}+\frac{12}{52}-\frac{6}{52}=\frac{26+12-6}{52}=\frac{32}{52}=\frac{8}{13}"


b)

"P=\frac{1}{4}+\frac{16}{52}-\frac{4}{52}=\frac{13+16-4}{52}=\frac{25}{52}"


c)

"P=\frac{1}{2}+\frac{1}{4}-\frac{1}{4}=\frac{1}{2}"